5. A 0.341-g sample of gas has a volume of 355 mL at 98.7 ℃ and 743 mmHg. What is the molar mass of this gas?
Answer #1

T = 98.7oC = 98.7oC + 273.15 = 371.85 K

P = 743 mmHg = 743 mmHg * (1 atm/760 mmHg) = 0.977 atm

V = 355 ml = 0.355 L

R = 0.0821 atm L mol-1 K-1

Ideal gas law: PV=nRT

n = PV/RT = (0.977 atm * 0.355 L)/(0.0821 atmLmol-1K-1 * 371.85 K) = 0.01136 mol

Given mass = 0.341 g

no. of moles, n = given mass/molar mass

=> molar mass = given mass/n = 0.341 g/0.01136 mol = 30.017 g/mol

Therefore, molar mass of the gas is 30.017g/mol.